A firm applies a production technology to produce units of output (q) according to the production function q = 20 x L^.5 x K^.5 , where L is a variable input and K is a fixed input. Assume K = 100 in the short run.

solve for the short-run average product of labor (AP L) and the marginal product of labor (MP L) functions. If L = 100, what are the numerical values of AP L and MP L?
Answer #1

q = 20L0.5K0.5  

K = 100

q = 20L0.5(100)0.5  

= 20(10)L0.5

q = 200L0.5

APL = q/L  

= 200L0.5/L  

= 200/L0.5   

L = 100

APL = 200/(100)0.5  

= 200/10  

= 20

q =   200L0.5

dq/dL= MPL = 200(0.5)L0.5 - 1

= 100L-0.5  

MPL = 100/L0.5  

L = 100

MPL = 100/(100)0.5

= 100/10

= 10